Bob invested $28,000, part at 19% and part at 17%. If the total interest at the end of the year is $4,920, how much did he invest at each rate?
Answer by elima(1433)

Bob invested $28,000, part at 19% and part at 17%. If the total interest at the end of the year is $4,920, how much did he invest at each rate?
x=amount at 19%
28000-x = amount at .17%
.19x+.17(28000-x)=4920
.19x+4760-.17x=4920
.19x-.17x=4920-4760
.02x=160
x=8000 -----amount at 19%
28000-(8000)=20000 -----amount at 17%
:)